Determine the pressure change when a constant volume of gas at 2.00 atm is heated from 30.0 °C to 40.0 °C.


1
Expert's answer
2021-05-13T07:21:39-0400

1"^{st}" condition:

"P_1=2atm"

"T_1=30^0C" "=30+273.15=303.15K"


2"^{nd}" condition:

Let "P_2=P_2"

"T_2=40^0C=40+273.15=313.15K"


Now from the gas equation we know that

"P\propto T"

Hence:


"\dfrac{P_1}{P_2}=\dfrac{T_1}{T_2}"


"P_2=2.067atm"


Now:


"\Delta P=P_2-P_1"

"\Delta P=2.067-2=0.067atm"